OpenStudy (gg):

Can somebody check this: solve differential equation y'y'''=2(y'')^2 if know that it's derivative of lower-ordered differential equation. I'll write my solution.

OpenStudy (nikvist):

Try, \[y'y'''=2(y'')^2\quad,\quad z=y'\]\[zz''=2(z')^2\]\[\frac{z''}{z'}=2\frac{z'}{z}\]\[(\ln{z'})'=2(\ln{z})'\quad\dots\]

OpenStudy (gg):

I did it in different way than wolfram did, but I didn't get the same result. I found a mistake :)